OpenStudy (anonymous):

To be able to do the problem 5(10 + 4) mentally, Zack does 50 + 20 instead of 5(14). Zack is using which of the following properties? associative commutative inverse distributive

OpenStudy (anonymous):

plz plz help

OpenStudy (texaschic101):

distributive property a(b + c) = ab + ac 5(10 + 4) = 5 * 10 + 5 * 4 = 50 + 20